An investment offers a 13.5 percent total return over the coming year. Bill Bernanke thinks the total real return on this investment will be only 4.5 percent. What does Bill believe the inflation rate will be over the next year? rev: 09_18_2012 Multiple Choice 8.61% -7.93% 9.90% 6.03% 11.20%
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.045)=(1+0.135)/(1+inflation rate)

1+inflation rate=(1.135/1.045)

inflation rate=(1.135/1.045)-1

=8.61%(Approx).